A semicircular arch over a street has a radius 10 feet. How high is the arch at a point whose ground distance is 4 feet from the center? Give the exact answer.
(please write answer as well tutors)

Answer by lwsshak3(11628)

A semicircular arch over a street has a radius 10 feet. How high is the arch at a point whose ground distance is 4 feet from the center? Give the exact answer.
***
Equation of circle with center at (0,0): x^2+y^2=r^2, r=radius
For given problem:
x^2+y^2=5^2
x^2+y^2=25
4 ft from center means x=4
4^2+y^2=25
y^2=25-16=9
y=±√9=±3
reject -3
How high is the arch at a point whose ground distance is 4 feet from the center? 3 ft